Mavis ate 3/10 of her birthday cake. She decided to give the brother 2/3 of what is left of the cake. What fraction of the cake was eaten by the brother?
Illustrate your answer through the use of diagrams,show the process clearly by making use of different types of shadings.
1
Expert's answer
2020-09-10T19:28:47-0400

Children have a cake(Picture 1).

Mavis ate 3/10 of her birthday cake. 

If they divided the cake into 10 pieces(Picture 2), 

Mavis ate 3 of them(Picture 3).

So there are 10 - 3 = 7 pieces of cake left(Picture 4) 

or 7/10 of a whole cake(Picture 5).

She gave 2/3 of what is left of the cake or 2/3 part of 7/10 of a whole cake(Picture 6).

It is like she divided the 7/10 of a whole cake into 3 pieces (Picture 7) and gave 2 of them to her brother(Picture 8).

Write a solution:

  1. 1-3/10 = 7/10 (left after Mavis ate part of the cake)
  2. 7/10 * 2/3 = 0.46(Mavis's brother ate)

Answer: 0.46 fraction of the cake was eaten by the brother(Picture 9)